# Pragmatic Reasoning improves LLM Code Generation

### v2

@@ 4 Experiment Setup
To analyze the merits of CodeRSA, we evaluate the performance of the three reranking methods (Coder, CoderReviewer, and CodeRSA) on a standard program code generation dataset using one of the latest language models. Since the strength of content-driven methods is their generality, we skipped extensive parameter tuning and simply used the default or commonly used parameters throughout our experiments.
4.1 Dataset and Model
The manually crafted HumanEval evaluation dataset (Chen et al., 2021) was designed to assess a model’s ability to convert natural language instructions into program code. It comprises 164 questions, each framed as an unfinished Python function starting with a brief instruction that describes how the function’s code should be completed.
We selected HumanEval for its proven role in code generation evaluation and balanced difficulty. For example, simpler datasets like CoNaLa (Yin et al., 2018) already yield near-perfect performance for the Coder model, leaving little room for reranking improvements. In contrast, more challenging datasets such as Bigcodebench (Zhuo et al., 2024) often produce many unsolvable instances, suggesting that more powerful models or improved problem modeling is needed.
Despite the availability of HumanEval for code generation, no dedicated and reliable benchmark currently exists for reranking generated code—a specialized subtask within code generation. Accordingly, the benchmark setup in our experiments involves three key stages: program code generation, testing (only for evaluation purposes), and reranking. Because LLM-based code generation is highly sensitive to hyperparameters and hardware configurations, we adopt a multi-sampling strategy. Specifically, for each of the 164 HumanEval problems, we first generate 300 candidate solutions at a temperature of 0.7. Then, we randomly select 50 problems, sample 10 candidates per problem, and repeat this process 10 times. Finally, we evaluate three different reranking methods on each resulting subset to measure their effectiveness.
In our experiments, we focus on the Llama-3-8B-Instruct variant, which comprises 8 billion parameters and is fine-tuned for instruction-following tasks Grattafiori et al. (2024). Llama-3-8B-Instruct is renowned for its lightweight design and robust performance, effectively balancing computational requirements with high-quality generation. Moreover, the HumanEval dataset poses a moderate challenge to the model, ensuring that it is neither overwhelmed by difficulty nor under-challenged. These characteristics render the model particularly suitable for evaluating the effectiveness of reranking methods in practical scenarios.
Additionally, we evaluated Llama-2-70b-chat Touvron et al. (2023) on a subset of HumanEval; however, its limited performance in code and instruction generation (and related poor probability estimates) led to its exclusion.
@@ 4.2 Implementation of Reranking Methods
Another baseline, “random”, is defined as the ratio of correct candidates to total candidates in each subset. By using the expected proportion of correct codes, this approach minimizes the impact of random seeds and naturally reflects the inherent difficulty of each sub-dataset.
State-of-the-art Method. Zhang et al. (2023) showed that CoderReviewer reranking (see Section 2 for details) outperforms Coder reranking and rivals execution-driven methods such as CodeT.
@@
In practice, we use the same prompt format as in Coder reranking to compute $\text{P}(c\mid i)$. To compute $\text{P}(i\mid c)$, the order of the instruction and the generated code snippet is reversed in the prompt (see Appendix A.5.2).
CodeRSA. To balance running time and limited computing resources, we restrict the RSA reranking procedure in our experiments to 10 candidates per question. We then apply a one-shot prompt (see Fig. 2, part B) to the LLM to generate one instruction for each candidate, yielding a set of potential instructions $\{i_{0},i_{1},\ldots,i_{10}\}$, where $i_{0}$ is the original instruction. Consistent with Coder reranking’s input prompt, we then compute the probability of each code candidate $c^{\prime}$ under each potential instruction $i^{\prime}$, forming an $11\times 10$ Coder score matrix. Then, we derive pragmatic speaker scores for each candidate with respect to the original instruction by contrasting how the candidate performs under $i_{0}$ versus under the other generated instructions. Finally, we use these pragmatic scores to rerank the candidates, selecting the one with the highest pragmatic speaker score for $i_{0}$.
@@ 5.1 Quantitative Analysis
In this section, we quantitatively evaluate three reranking methods, including CodeRSA, in terms of accuracy using ten subsets sampled from the HumanEval dataset and their corresponding generated code candidates.
Figure 3: Box plot of accuracy for the reranking methods. The orange line indicates the mean accuracy.
Here, accuracy is defined as the fraction of candidates selected by a reranking method that pass all test cases relative to the total number of candidates in the subset. Fig. 3 presents box plots comparing the accuracy distributions of the three reranking methods and a random baseline. Variance is estimated from bootstrapping across ten subsets. CodeRSA attains the highest mean accuracy, with relatively narrow interquartile ranges indicating consistent performance. CoderReviewer shows moderate accuracy, with a mean exceeding Coder’s but still below CodeRSA. Coder reranking exhibits a relatively large variance, indicating that its performance is less stable and more prone to fluctuations, likely due to its tendency to favor overly generic solutions.
@@
Figure 4: Two code candidates along with relevant information
Coder’s lower performance suggests that relying solely on a literal listener perspective is insufficient for effective reranking and that CodeRSA’s more complex reasoning leads to superior performance, demonstrating the benefits of a comprehensive RSA modeling approach.
@@ 5.2 Qualitative Analysis
Zhang et al. (2023) observed that language models often generate “degenerate” programs—overly generic yet repetitive code candidates. In our analysis, we observe that Coder reranking indeed favors these degenerate outputs because it evaluates the cumulative token likelihood of a candidate $c$, given the original instruction $i_{0}$:
where $c^{(t)}$ denotes the token at position $t$, and $c^{(<t)}$ represents the sequence of the preceding tokens. Each factor $\text{P}_{\text{LLM}}(c^{(t)}\mid i_{0},c^{(<t)})$ is strictly less than 1; thus, longer sequences accumulate a lower overall probability, biasing Coder reranking towards shorter, potentially degenerate code. Nonetheless, differences in token-level probabilities can offset this bias: a longer but logically correct candidate may achieve a higher overall probability if each token is assigned a sufficiently high probability.
With recent advances in LLMs, generating degenerate programs has become less common, although incomplete functionality remains a concern. For example, in Fig. 4, the original instruction requires returning the greatest integer above zero whose frequency is, at least, its own value, or $-1$ if none exists. Nevertheless, code_09 omits the “greatest” integer requirement and the $-1$ fallback, making it shorter and more likely to be highly ranked by Coder.
Fig. 4 presents the Coder scores $\log\text{P}_{L_{0}}(c\mid i)$ for function_01 and function_09 under all potential instructions (including $i_{0}$). Additionally, code_01 achieves a log-probability of $-29.24$ for $i_{0}$, whereas code_09 reaches $-21.12$. Consequently, Coder reranking selects the incomplete code_09. The Reviewer component alone is insufficient to offset code_09’s higher Coder score; hence CoderReviewer also prefers code_09. The difference between Coder and Reviewer scores highlights a major flaw in the CoderReviewer method. Since these scores often differ significantly, treating them as equally important is not optimal. Additionally, trying to adjust their weights can add more complexity and uncertainty.
Fig. 4 reports the pragmatic speaker scores $\text{R}_{S_{1}}(i\mid c)$ for each instruction–code pair. Notably, code_01 receives a score of $-0.0458$ with $i_{0}$, whereas code_09 has $-0.0533$. Acting as a pragmatic listener, CodeRSA selects code_01, which more closely aligns with $i_{0}$ from a pragmatic speaker’s perspective.
Although the score calculation in our method uses log probabilities, two observations confirm that our method still achieves the core intuitions put forward by the RSA framework:
Firstly, as can be seen in Fig. 4, code_01 has a lower overall Coder score than code_09, likely because it is longer and more complicated. However, under $i_{0}$, code_01 obtains a higher Coder score relative to other possible instructions, which then translates into a higher pragmatic speaker score after normalization. This process shows that CodeRSA examines each candidate’s relative scores across all potential instructions to verify whether it truly ranks highly under the original instruction.
Secondly, as shown in Fig. 4, code_09 receives multiple high Coder scores, especially under the instruction generated by itself. From the literal listener’s perspective, the self-generated instruction more precisely describes code_09, raising its overall Coder scores. Crucially, because the speaker’s score is normalized by the sum of log probabilities across all instructions for a given candidate, a higher Coder score on the self-generated instruction will reduce the relative weight assigned to the original instruction. Consequently, the pragmatic speaker will prefer to use another instruction for code_09 instead of the user’s original instruction. Moreover, as can be seen in Fig. 4, the self-generated instruction for code_09 matches its details more closely.
@@ 6 Discussion
Our proposed CodeRSA approach contains a number of simplifications compared to the original RSA model, which has been developed for describing human-human communication: (1) we assumed a uniform speaker cost for the instructions. While this simplification makes the analysis more tractable, it means that our model does not currently take into account effects related to how “costly” an instruction would be to produce for the human speaker. Future work should investigate variable cost structures to better capture these nuances.
We also assume that all code candidates are equally likely a priori, effectively omitting the candidate prior $\text{P}(c)$ from our calculations. Although this assumption allows us to concentrate solely on the pragmatic speaker’s score, it might not reflect real-world scenarios where some code candidates are more common due to usage frequency, domain-specific patterns, or contextual relevance. Future work should explore the use of non-uniform priors to better capture candidate plausibility.
In section 4, we argued that our approach, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates. This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication. Research on human communication has demonstrated the importance of pragmatic reasoning in communication despite its apparent computational overhead, while at the same time suggesting that humans may also learn to use simple heuristics or amortized estimates (Pu et al., 2024) to not engage in iterative reasoning in easy cases, while still being able to employ the full reasoning procedure in more difficult cases where it is necessary.
@@ 8 Limitations
A known limitation of RSA approaches is their computational complexity and associated resource consumption. For example, on a single NVIDIA Tesla A100 (PCIe 4.0, 80GB HBM2e, 300W), performing complete CodeRSA inference on 500 instances takes nearly 8 hours. Our approach compares each potential instruction with every candidate, leading to a quadratic increase in complexity as the number of candidates grows. Although CodeRSA can theoretically handle many candidates, we limited our experiments to ten candidates per question to keep runtime and hardware usage manageable. This restriction inevitably narrows the variety of solutions and may affect how well the approach generalizes to larger-scale scenarios.
@@
Another limitation is that while we employed a multi-sampling strategy to mitigate uncertainty, we have only utilized a single dataset and one model so far. We are currently working on incorporating several balanced-difficulty datasets—such as MBPP (Austin et al., 2021) and DS-1000 (Lai et al., 2023)—along with additional open-source models like Mistral (Jiang et al., 2023) and Qwen (Bai et al., 2023). This expansion will allow us to further assess the effectiveness of different reranking methods across diverse scenarios, ultimately leading to a more robust evaluation of our approach.
@@ A.2 More Details of Results
Table 1 summarizes the average accuracy and standard deviation across ten sampled subsets.
Table 1: Mean accuracy (with standard deviation) across 10 subsamples for each reranking method.
Method Mean Accuracy Std. Dev.
Coder 0.48 0.059
CoderReviewer 0.51 0.043
CodeRSA 0.55 0.044
random 0.43 0.030
As shown in Table 1, CodeRSA achieves the highest average accuracy (0.55), outperforming Coder (0.48), CoderReviewer (0.51), and the random baseline (0.43). This indicates that CodeRSA provides stronger overall performance compared to the other methods. From a standard-deviation perspective, CoderReviewer and CodeRSA both exhibit relatively stable performance across different subsamples. By contrast, Coder has the highest standard deviation (0.059), suggesting greater sensitivity to varying combinations of coding problems.
Figure 5: Performance difference between CodeRSA and CoderReviewer.
Fig. 5 depicts the per-trial accuracy difference between CodeRSA and CoderReviewer (i.e., CodeRSA accuracy minus CoderReviewer accuracy). Bars above zero indicate that CodeRSA outperforms CoderReviewer, while the single negative bar (Trial 2) represents the only instance where CodeRSA yields a lower accuracy. In most trials, CodeRSA demonstrates gains of up to 10%, highlighting its consistent advantage.
A.3 Ablation Study
Figure 6: Box plot of accuracy in ablation study.
To investigate the impact of the log transformation in the calculation of the pragmatic speaker score, we conducted an ablation study comparing the CodeRSA’s modeling of pragmatic speaker score:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
against a variant that directly utilizes the normalized probability:
$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
referred to as CodeRSA_nolog.
Fig. 6 displays the box plots of accuracy for different methods on ten subsets. From the results, CodeRSA achieves the highest mean accuracy with a narrow interquartile range, indicating stable performance. In contrast, CodeRSA_nolog shows a marked decrease in accuracy, evidenced by a lowest mean and a compressed interquartile range. This observation aligns with previous findings by Schuster et al. (2024), which reported that the basic RSA model underperforms the baseline on their SpreadNaLa dataset. These results suggest that applying a logarithmic transformation to expand the probability intervals can substantially enhance overall performance.
A.4 A Geometric Mean Perspective on the Log-Normalization Formula
In our proposed CodeRSA model, we define the pragmatic speaker score as follows:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
where $\text{P}_{L_{0}}(c\mid i)$ is the probability estimated by the literal listener for candidate code $c$ given instruction $i$, and $I$ denotes the set of potential instructions. The geometric mean of the set $\{\text{P}_{L_{0}}(c\mid i^{\prime}):i^{\prime}\in I\}$ is defined as:
$\text{GM}=\left(\prod_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})\right)^{\frac{1}{|I|}},$
and its logarithm is given by:
$\log\text{GM}=\frac{1}{|I|}\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime}).$
Thus:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{|I|\cdot\log\text{GM}}\;.$
In this sense, normalizing the log probability for a particular instruction $i$ by the sum over all potential instructions can be seen as a way of capturing its relative contribution in a manner analogous to the geometric mean.

### v3

@@ 4 Experiment Setup
To understand the strengths and weaknesses of CodeRSA, we evaluate the performance of three reranking methods (Coder, CoderReviewer, and CodeRSA) on widely used benchmarks for code generation. Since the advantage of content-driven methods lies in their generality, we rely on commonly adopted default settings and perform only minimal sensitivity checks on key parameters.
4.1 Dataset and Base Models
We evaluate on two widely used code generation benchmarks. HumanEval (Chen et al., 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction. MBPP (Austin et al., 2021) includes 257 short programming tasks with natural language prompts. HumanEval offers balanced difficulty, while MBPP introduces greater lexical variety. Note that simpler datasets such as CoNaLa (Yin et al., 2018) already yield near-perfect results, leaving little room for reranking, whereas more challenging datasets such as BigCodeBench (Zhuo et al., 2024) contain many instances that cannot yet be solved by today’s state-of-the-art models, which makes it difficult to obtain meaningful comparisons of reranking methods and may obscure the performance differences we aim to study.
We use the following setup: for each problem in HumanEval and MBPP, we sample $n=10$ candidate codes at a temperature of 1.0. We then evaluate reranking methods on this shared candidate set. A sensitivity check with varying numbers of sampled candidates is provided in Appendix A.3.
For our experiments, we use Llama-3-8B-Instruct (Grattafiori et al., 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale. Llama-3-8B-Instruct balances efficiency with strong generation quality, while Qwen-2.5-7B-Instruct provides a competitive open-source alternative. Both achieve competitive performance on HumanEval and MBPP, making them suitable for assessing reranking in our setting. We do not include specialized coder models in this study, since our framework requires both code generation and instruction-level reasoning. Future work could explore hybrid setups, for example, using coder models for program synthesis combined with general-purpose instruction-tuned models for reasoning about instructions.
@@ 4.2 Implementation of Reranking Methods
State-of-the-art Method. Zhang et al. (2023a) showed that CoderReviewer Reranking (see Section 2 for details) outperforms Coder Reranking and rivals execution-driven methods such as CodeT. In practice, we use the same prompt format as in Coder Reranking to compute $\text{P}(c\mid i)$. To compute $\text{P}(i\mid c)$, the order of the instruction and the generated code snippet is reversed in the prompt (see Appendix A.5.2).
@@
CodeRSA. To balance runtime and computational constraints, we limit the process to $n=10$ candidate programs per problem. For CodeRSA, we further generate one additional instruction ($m=1$) for each candidate using a one-shot prompt (temperature = 0.7; see Fig. 2, part B).
Rather than treating each instruction independently, we next use the same LLM that generated the candidates to perform pairwise semantic equivalence judgments among these instructions. Following prior work on LLM-based clustering (Zhang et al., 2023b), we cluster instructions that express the same functionality (see Section 3). Concretely, we query the LLM with a 3-shot prompt containing examples of both positive and negative semantic equivalence pairs, asking it to judge whether two instructions express the same functionality. We then build a pairwise equivalence graph where each node represents an instruction and an edge indicates semantic equivalence according to the LLM. The connected components of this graph are treated as clusters of mutually equivalent instructions (see Appendix A.4).
For each candidate $c$, we compute literal listener scores with respect to every instruction, and then aggregate them at the cluster level: non-main clusters take the mean across their members, while the cluster containing the original instruction $i_{0}$ retains its direct probability. We also incorporate candidate priors through a candidate-specific temperature parameter $\tau_{c}$, controlled by a coefficient $\alpha$ (see Section 3 for the definition). Finally, we apply softmax normalization with these temperatures over clusters to obtain cluster-level pragmatic speaker scores. The pragmatic listener then reranks candidates by selecting the one with the highest speaker score with respect to the $i_{0}$-cluster, which represents the original user intent.
@@ 5.1 Quantitative Analysis
In this section, we analyze the quantitative performance of CodeRSA with respect to the calibration parameter $\alpha$ (defined in Section 3), which controls the influence of the prior through temperature scaling, using the MBPP dataset and Llama-3-8B-Instruct model.
Figure 3: Accuracy of CodeRSA across different values of the calibration parameter $\alpha$. The shaded region indicates a stable performance band.
Here, accuracy denotes the proportion of test instances where the reranking method selects a candidate that passes all test cases provided by the benchmark. Figure 3 shows that CodeRSA consistently outperforms baseline reranking methods, with clustering playing a crucial role: removing the clustering step yields a substantial drop in accuracy. The figure also highlights the robustness of CodeRSA to the choice of $\alpha$: within the stable band of $[0.90,1.15]$, performance remains consistently above both Coder and CoderReviewer. At $\alpha=1.0$, CodeRSA achieves the best accuracy of 59.53%, clearly surpassing the baselines. These results demonstrate that CodeRSA’s pragmatic reasoning, enhanced by clustering, is not overly sensitive to calibration.
We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same overall trends.
@@
Figure 4: Qualitative Example: Bias in Coder vs. CodeRSA Correction
@@ 5.2 Qualitative Analysis
Zhang et al. (2023a) note that reranking based on cumulative token likelihood tends to prefer shorter candidates, since each token probability is <1<1 and longer sequences accumulate lower overall scores. This bias makes the Coder approach prone to favoring incomplete or generic programs. In Fig. 4, the instruction requires returning the greatest integer above zero whose frequency is at least its own value, or 1-1 if none exists. However, code_09 omits both the “greatest” requirement and the 1-1 fallback, making it incomplete but shorter. As shown in Fig. 4, Coder assigns code_09 a higher score (21.12-21.12) than the correct code_01 (29.24-29.24), and thus prefers the degenerate solution. CoderReviewer inherits this issue, as Reviewer alone cannot offset code_09’s inflated Coder score.
Fig. 4 reports CodeRSA’s cluster-level pragmatic speaker scores after softmax normalization ($\alpha=1$). Instructions with equivalent semantics are grouped, and the cluster containing the original instruction $i_{0}$ is treated as the main cluster. Here, code_01 achieves a score of 0.32 on the main cluster, compared to 0.25 for code_09. Notably, code_09 also receives relatively high confidence on Cluster_4, which dilutes its probability on the main cluster due to RSA normalization. In RSA terms, code_09 is not strongly aligned with either the main cluster or Cluster_4, indicating that it fits the intended instruction less well than other candidates. By contrast, the probability of code_01 is concentrated on the main cluster, which better aligns with the original instruction and is therefore favored under pragmatic reasoning.
Taken together, this case study shows how CodeRSA operationalizes RSA reasoning: By normalizing over alternative clusters, it penalizes candidates that spread probability mass across multiple interpretations and favors those that focus on the main cluster, thereby improving robustness and faithfulness in reranking.
@@ 6 Discussion
Our proposed CodeRSA approach contains a number of simplifications compared to the original RSA model, which has been developed for describing human–human communication: It assumes a uniform speaker cost for the instructions. While this simplification makes the analysis more tractable, it means that our model does not currently take into account effects related to how “costly” an instruction would be to produce for the human speaker. Future work should investigate variable cost structures to better capture these nuances.
In Section 4, we argued that CodeRSA, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates. This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication. Research on human communication has demonstrated the importance of pragmatic reasoning, even though it introduces additional computational overhead. At the same time, studies suggest that humans may rely on simple heuristics or amortized estimates (Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex.
@@ 8 Limitations
A known limitation of RSA approaches is their computational complexity and associated resource consumption. For example, on a single NVIDIA Tesla A100 (PCIe 4.0, 80GB HBM2e, 300W), performing complete CodeRSA inference on 500 instances takes nearly 6 hours. Our approach compares each potential instruction with every candidate, leading to a quadratic increase in complexity as the number of candidates grows. Although CodeRSA can theoretically handle many candidates, we limited our experiments to ten candidates per question to keep runtime and hardware usage manageable. This restriction inevitably narrows the variety of solutions and may affect how well the approach generalizes to larger-scale scenarios.
@@
Another limitation is that, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the scope remains relatively narrow. We are currently working on incorporating additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023)), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5. This expansion will allow us to evaluate reranking methods across a wider range of scenarios, ultimately leading to a more comprehensive assessment of our approach.
@@ A.2 More Details of Results
In this subsection, we provide further results on different datasets and models to further validate the robustness of CodeRSA. Specifically, we evaluate on the HumanEval dataset as well as on MBPP, using both Llama-3-8B-Instruct and Qwen2.5-7B-Instruct.
Figure 5: Accuracy of CodeRSA on HumanEval with Qwen2.5-7B-Instruct. The shaded region shows the stable band.
Figure 6: Accuracy of CodeRSA on HumanEval with Llama-3-8B-Instruct. The shaded region shows the stable band.
Figure 7: Accuracy of CodeRSA on MBPP with Qwen2.5-7B-Instruct. The shaded region shows the stable band.
Across all settings, several consistent trends can be observed.
Clustering effectiveness. Removing the clustering step leads to a noticeable drop in accuracy, highlighting its role in reducing redundancy and stabilizing pragmatic reasoning.
Calibration robustness. CodeRSA is not strongly sensitive to the calibration parameter $\alpha$; performance remains stable across a relatively wide range rather than relying on a finely tuned value.
Superior accuracy. CodeRSA consistently achieves higher accuracy than both baselines. On the HumanEval dataset, we observe some fluctuations in performance, and the overall accuracy is relatively low. This may be partly due to randomness or parameter settings in the experiments. However, since all reranking methods are evaluated under the same conditions, the relative comparison between them remains fair and informative.
These findings confirm that the improvements achieved by CodeRSA are reliable across different models and datasets. The calibration parameter $\alpha$ is shown to be both interpretable and stable, further supporting the practicality of the approach.
A.3 Accuracy vs. Number of Sampled Candidates (MBPP)
Figure 8: MBPP: Mean accuracy vs. number of randomly sampled candidates per problem ($n$). Curves show CodeRSA (orange), Coder (green), and CoderReviewer (purple). Shaded regions indicate the standard deviation across multiple random samplings.
To further examine the impact of candidate diversity on reranking performance, we conducted a controlled study varying the number of sampled code candidates per MBPP problem ($n=1\ldots 10$). For each value of $n$, we randomly sampled $n$ candidates from the pool of ten generated solutions and applied three reranking strategies: Coder, CoderReviewer, and our proposed CodeRSA. The experiment was repeated ten times with different random seeds for each value of $n$, and the figure reports the mean accuracy and its standard deviation across runs.
When $n=1$, all methods yield identical results, as no reranking can occur. As $n$ increases, performance improves for all methods due to a broader candidate set, but the gain plateaus after approximately $n=7$. Across all sampling levels, CodeRSA achieves the highest accuracy, maintaining a margin of roughly 2–3 percentage points over CoderReviewer and up to 5 points over Coder. This shows that pragmatic reasoning allows CodeRSA to better leverage candidate diversity while remaining robust to sampling variability. The narrow confidence bands further indicate stable performance even under random candidate selection, confirming its reliability when generation stochasticity varies across runs or models.
In our main experiments, we fixed $n=10$ candidates per problem as a practical balance between computational cost and runtime. The results here further suggest that model performance is not strongly dependent on candidate set size. Future work could explore larger candidate pools when computational resources permit.
A.4 An example of clustering
In the following presentation, each item is denoted in the format:
$\textbf{code\_X}:\text{Instruction generated from this code}$
This means that the left-hand side (code_X) represents the identifier of the function implementation, and the right-hand side is the instruction generated based on it.
Main Cluster: Maximum value with frequency condition
  • code_1: Create a function that takes a list of numbers. Returns the largest number that appears at least as many times as its value in the list. If no such number exists, return -1.
  • code_6: Create a function that takes a list of integers and returns the maximum value that appears at least as many times as its value. If no such value exists, return -1.
  • code_8: Create a function that takes a list of numbers and returns the maximum integer that occurs at least as many times as its value. If multiple such numbers exist, return the largest one. If no such number exists, return -1.
Cluster 2: Most frequent element
  • code_5: Create a function that takes a list of integers. Returns the number that appears most frequently in the list. If there are multiple such numbers with the same frequency, return the largest one.
  • code_10: Create a function that takes a list of integers. Returns the most frequent integer greater than 0. If multiple integers have the same highest frequency, return the smallest one. If the list is empty, return -1.
Cluster 3: Repeated integers
  • code_4: Create a function that takes a list of integers and returns the smallest positive integer that appears more than once. If no such integer exists, return -1.
  • code_7: Create a function that takes a list of integers and returns the maximum value that appears more than once. If no such value exists, return -1.
Cluster 4: Missing positive integer
  • code_3: Create a function that takes a list of integers and returns the first missing positive integer. If the list is empty, return -1.
Cluster 5: First/last integer with frequency condition
  • code_2: Create a function that takes a list of integers and finds the first integer that occurs at least as many times as its value. If no such integer is found, return None.
  • code_9: Create a function that takes a list of integers. Returns the first integer that appears at least as many times as its value in the list. If no such integer exists, return None.